OpenStudy (anonymous):

6x+ 1/2y = 3 standard form to slope- intercept form ?

OpenStudy (anonymous):

@emilyxo follow as done earlier

OpenStudy (anonymous):

What you have to do is try to get the y alone. Subtract 6x from both sides. That gives you 1/2y=-6x+3 Now, you have to make your y a whole number. To do this, you multiply both sides by 2/1. Your answer is y= -12x +6

OpenStudy (anonymous):

6x+ 1/2y = 3 multiplying by 2 both sides we have 12x+y=6 y=-12x+6 here m=-12 and c=6

OpenStudy (anonymous):

thanks :)

OpenStudy (aravindg):

slope intercept form is y=mx+c